# GLANCE: Graph Logic Attention Network with Cluster Enhancement for Heterophilous Graph Representation Learning

### v1

@@ Abstract
Keywords:
Graph Representation Learning Logic Networks Heterophily
@@ 1 Introduction
Graph Neural Networks (GNNs) have become a pivotal tool for analyzing graph-structured data, with applications spanning recommendation systems,finance, biomedicine, and social networks [16, 2, 12, 10]. These models excel in scenarios characterized by homophily, where connected nodes share similar features or labels [4]. However, many real-world graphs, such as protein interaction networks and citation graphs, exhibit heterophily, where connected nodes differ significantly in their attributes or class labels. Traditional GNNs often struggle in such settings, as their message-passing mechanisms can propagate irrelevant or misleading information, diminishing their effectiveness [14, 17, 11].
@@ 3.4 Dynamic Graph Refinement
[…]

### v2

@@ 1 Introduction
Graph Neural Networks (GNNs) have become a pivotal tool for analyzing graph-structured data, with applications spanning recommendation systems,finance, biomedicine, education and social networks [25, 17, 18, 5, 6, 15, 21, 2, 19, 14]. These models excel in scenarios characterized by homophily, where connected nodes share similar features or labels [8]. However, many real-world graphs, such as protein interaction networks and citation graphs, exhibit heterophily, where connected nodes differ significantly in their attributes or class labels. Traditional GNNs often struggle in such settings, as their message-passing mechanisms can propagate irrelevant or misleading information, diminishing their effectiveness [23, 26, 16, 20, 4, 3].
@@ 3.4 Dynamic Graph Refinement
where $H$ is the number of attention heads, $\mathbf{w}_{h}$ is the weight vector for head $h$, $\|$ denotes concatenation, and $\sigma$ is the sigmoid function.